# Optimal Rank-Metric Codes with Rank-Locality from Drinfeld Modules

### v1

@@ 2.1. Drinfeld Modules
See [19, Theorem 7.1.19]
@@
Let $R,\ell\geq 1$. Let $\{a_{1},\ldots,a_{\ell}\}\subseteq{\mathbb{F}}_{q}^{*}$. Let $h=\prod_{i=1}^{\ell}(T^{R}-a_{i})$ and $P\equiv 1\mod h$ be a monic irreducible polynomial in ${\mathbb{F}}_{q}[T]$ with residue field ${\mathbb{F}}_{q^{m}}$. Let $\phi^{(1)}$ be the Drinfeld module of rank $1$ defined in Theorem 2.12 with $g=1$ and $\phi_{T}^{(1)}=T+\tau$. Then, the $h$-division polynomial of $\bar{\phi}^{(1)}$ splits completely in ${\mathbb{F}}_{q^{m}}$.
@@
Observe that $g=1$, and from the choice of $P$ we get immediately that $P$ is a totally split place for the $h$-division field $M={\mathbb{F}}_{q}(T)(\phi^{(1)}[h])$. This means that all places of ${\mathbb{F}}_{q}(T)(\phi^{(1)}[h])$ lying above $P$ have relative degree $1$ over the residue field at $P$. In other words, if $\alpha\in{\mathbb{F}}_{q^{m}}$ is a root of $P$, and $R$ is any place above $P$, then $[\mathcal{O}_{R}/R:{\mathbb{F}}_{q}(\alpha)]=1$. Now we want to see that this leads to a complete split of $\bar{\phi}^{(1)}_{h}(x)$ in ${\mathbb{F}}_{q}(\alpha)[x]\cong{\mathbb{F}}_{q^{m}}[x]$. Now let $\beta$ be a root of $\phi^{(1)}_{h}(x)$ in $M$ and consider the intermediate extension $L={\mathbb{F}}_{q}(T)(\beta)$. Then we have $M\supseteq L\supseteq{\mathbb{F}}_{q}(T)$. If $P$ splits completely in $M:{\mathbb{F}}_{q}(T)$, then $P$ splits completely in $L:{\mathbb{F}}_{q}(T)$ since ramification and relative degrees are multiplicative. So we have that $\bar{\phi}^{(1)}_{h}(x)\mod P$ cannot have irreducible factors of degree larger than $1$ because of Kummer Theorem [24, Theorem 3.3.7]. It remains to show that no root of $\bar{\phi}^{(1)}_{h}(x)\mod P$ has multiplicity larger than $1$, so that the number of irreducible factors of $\bar{\phi}^{(1)}_{h}\mod P$ is exactly $\deg(\phi^{(1)}_{h}(x))=q^{R\ell}$, which forces that $\phi^{(1)}_{h}(x)$ to be totally split, as wanted. For $\beta\in\overline{\mathbb{F}}_{q}$ let us denote by $m_{\beta}$ the minimal polynomial of $\beta$ over ${\mathbb{F}}_{q}$. To see that $\bar{\phi}^{(1)}_{h}\mod P$ has no repeated roots, observe that the only specializations $\beta\in\overline{\mathbb{F}}_{q}$ of $T$ which lead to a repeated root of $\bar{\phi}^{(1)}_{h}(x)\mod m_{\beta}$ are in correspondence with the zeroes in $\overline{{\mathbb{F}}}_{q}$ of the coefficient of $x$ in $\phi^{(1)}_{h}(x)$ (simply because $\phi^{(1)}_{h}(x)$ is $q$-linearized). But, by direct computation, one sees that this coefficient is $h(T)$, and $P$ is coprime with $h$. This forces that $\alpha$ is never a specialization that leads to a repeated root of $\phi^{(1)}_{h}(x)$ or, more precisely, $\bar{\phi}^{(1)}_{h}\mod P$ is squarefree, and therefore totally split (since each of its irreducible factors has degree $1$). ∎
Observe that $K\{\tau\}$ can be seen as module over $K[T]$. For any $n\in{\mathbb{N}}$, we denote by $K\{\tau\}_{\leq n}$ the set $\{\sum^{n}_{i=0}a_{i}\tau^{i}\;|\;a_{i}\in K\}$.
Proposition 2.14.
$K\{\tau\}$ is free module of rank $R$ over $K[T]$.

### v2

@@ 3. General Theorem and Main Construction
Since $\gcd(\mathfrak{p},h)=1$, by Proposition 2.10, we have
$\operatorname{ht}(\phi_{h})=H(\phi)\cdot\operatorname{ord}_{\mathfrak{p}}(h)\cdot\deg(\mathfrak{p})=0.$
This means that $\phi_{h}$ is separable, and since $\deg_{\tau}(\phi_{h})=\ell R$, then $\dim_{{\mathbb{F}}_{q}}\phi[h]=\dim_{{\mathbb{F}}_{q}}\operatorname{Ker}(\phi_{h}(x))=\ell R$. In particular, since $\gcd(\mathfrak{p},h)=1$, it is also $\gcd(\mathfrak{p},T-a_{i})=1$ for every $i=1,\ldots,\ell$. Therefore, applying the same argument we prove our claim. ∎
3.1. Main Code Construction
In this subsection we prove Theorem 3.1. Let us start the construction of our code by choosing distinct $a_{1},\dots a_{\ell}\in{\mathbb{F}}_{q}^{\times}$ and $h=\prod^{\ell}_{i=1}(T-a_{i})$ for some $\ell>0$. Let $\phi$ be a Drinfeld module of rank $R=r+\delta-1$ over ${\mathbb{F}}_{q^{m}}$ with ${\mathbb{F}}_{q}[T]$-characteristic being a monic irreducible polynomial $\mathfrak{p}$ of degree $m$ such that $\gcd(\mathfrak{p},h)=1$ and ${\mathbb{F}}_{q^{m}}(\phi[h])\subseteq K={\mathbb{F}}_{q^{tm}}$, for some $t\geq 1$.
The following conditions are a consequence of the hypothesis of Theorem 3.1. We will be using them to prove the theorem.
  1. (C0)
    $\ell<q$: since the $a_{i}$’s have to be distinct,
  2. (C1)
    $s+1\leq\ell$: included in the hypothesis,
  3. (C2)
    $\ell R\leq tm$: notice that, since $h$ has all distinct roots and $\gcd(\mathfrak{p},h)=1$, then by Proposition 2.10 we have that $\phi_{h}$ has all distinct roots from which it follows $\dim_{{\mathbb{F}}_{q}}(\phi[h])=\ell R\leq\dim_{{\mathbb{F}}_{q}}(K)$.
Roughly, these conditions simply allow for enough space in ${\mathbb{F}}_{q}^{tm\times\ell R}$ to construct a code with the wanted dimension and $(r,\delta)$-rank-locality.
The message space $\mathcal{M}$ is given by
$\mathcal{M}=\left\{\sum^{s}_{k=0}g_{k}(\tau){\phi}_{T}^{k}\mid g_{k}\in{\mathbb{F}}_{q^{tm}}\{\tau\}_{\leq r-1}\right\}.$
For $i=1,\ldots,\ell$ let $W_{i}=\phi[T-a_{i}]$ and observe that by Propostion 2.14 we have that $\phi[h]=\oplus^{\ell}_{i=1}W_{i}$. Let $\{\beta^{(i)}_{1},\dots,\beta^{(i)}_{R}\}$ be a basis for $W_{i}$. The encoding map is then given by
$\operatorname{enc}:\mathcal{M}$ $\longrightarrow{\mathbb{F}}^{\ell R}_{q^{tm}}$
C1 guarantees that the encoding is injective, and therefore one obtains the wanted dimension. Moreover, the fact that ${\mathbb{F}}_{q^{m}}(\phi[h])\subseteq K$ gives that the $\beta_{j}^{(i)}$’s are independent elements of $K={\mathbb{F}}_{q^{tm}}$.
Theorem 3.3.
The code defined by $\operatorname{enc}(\mathcal{M})$ is an optimal
$\left(tm\times\ell R,tm(s+1)r,\ell R-Rs-r+1,r,\delta\right)$
(matrix) rank-metric code w.r.t. the bound in Proposition 2.22.
Proof.
Let $W=\phi[h]=\phi\left[\prod_{i=1}^{\ell}(T-a_{i})\right]$. To prove the length of the code is $\ell R$, first, we show that the $W_{i}$’s are orthogonal. Observe that $\phi[T-a_{i}]$ is an ${\mathbb{F}}_{q}[T]$-module. Given $\operatorname{gcd}(T-a_{1},\ldots,T-a_{\ell})=1$ then, by Proposition 2.14, we have
$W=\phi\left[\prod_{i=1}^{\ell}(T-a_{i})\right]=\prod_{i=1}^{\ell}\phi[T-a_{i}]=\prod_{i=1}^{\ell}W_{i}.$
Now, if we consider $W_{i}$’s as ${\mathbb{F}}_{q}$-vector spaces, since $W_{j}\cap\bigoplus_{\scriptstyle\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{\ell}W_{i}=\{0\}$, then $W=\bigoplus_{i=1}^{\ell}W_{i}$. In particular, $\{\beta_{j}^{(i)}\}_{\begin{subarray}{c}1\leq i\leq\ell\\ 1\leq j\leq R\end{subarray}}\subseteq{\mathbb{F}}_{q^{tm}}$ is an ${\mathbb{F}}_{q}$-basis of $W$ and thus $\dim_{{\mathbb{F}}_{q}}(W)=|\{\beta_{j}^{(i)}\}_{\begin{subarray}{c}1\leq i\leq\ell\\ 1\leq j\leq R\end{subarray}}|=\ell R$. This shows that the length of $\operatorname{enc}(\mathcal{M})$ (as a vector rank-metric code) is $\ell R$. Finally, by fixing an ${\mathbb{F}}_{q}$-basis for ${\mathbb{F}}_{q^{tm}}$, each element in $\operatorname{enc}(\mathcal{M})$ can be seen as a matrix in ${\mathbb{F}}_{q}^{tm\times\ell R}$.
We now show that the ${\mathbb{F}}_{q^{tm}}$-dimension of the code is $(s+1)r$. Set $f(\tau)\in\mathcal{M}$. From the construction, $f(\tau)=\sum^{s}_{k=0}g_{k}(\tau){\overline{\phi}}_{T}^{k}$ for some $g_{k}\in{\mathbb{F}}_{q^{tm}}\{\tau\}_{\leq r-1}$. Now, for every $k\in\{0,\ldots,s\}$, we have $g_{k}(\tau)=\sum_{j=0}^{r-1}a^{(k)}_{j}\tau^{j}$ where $a^{(k)}_{j}\in{\mathbb{F}}_{q^{tm}}$ for every $j\in\{0,\ldots,r-1\}$. This means that every vector
$\bm{a}=\left(a^{(0)}_{0},\ldots,a^{(0)}_{r-1},\ldots,a^{(k)}_{0},\ldots,a^{(k)}_{r-1},\ldots,a^{(s)}_{0},\ldots,a^{(s)}_{r-1}\right)\in{\mathbb{F}}_{q^{tm}}^{(s+1)r}$
can be associated to $g_{0}(\tau),\ldots,g_{s}(\tau)\in{\mathbb{F}}_{q^{tm}}\{\tau\}_{\leq r-1}$, and thus to each $f(\tau)\in\mathcal{M}$. This shows that the ${\mathbb{F}}_{q^{tm}}$-dimension of the code is $(s+1)r$ as long as the encoding map is injective. To see the injectivity of the map $\operatorname{enc}$, simply observe that $\deg_{\tau}(\phi_{T})=R$, and a non-zero polynomial of linearized degree $sR+r-1<(s+1)R\leq\ell R$ cannot be the zero map on a space of ${\mathbb{F}}_{q^{tm}}$-dimension $\ell R$. This shows that the ${\mathbb{F}}_{q}$-dimension of the code is $tm(s+1)r$.
We now need to establish $(r,\delta)$-rank-locality according to Definition 2.21. Considering the code as a matrix rank-metric code (i.e. codewords are matrices), suppose $\gamma\in\{1,\ldots,\ell R\}$ is a column index. Then, the $\gamma$-th column in every codeword is an entry of the form $f(\beta^{(i_{\gamma})}_{j_{\gamma}})$ for some $i_{\gamma}\in\{1,\ldots,\ell\}$ and $j_{\gamma}\in\{1,\ldots,R\}$ where $\beta^{(i_{\gamma})}_{j_{\gamma}}$ is an element of the ${\mathbb{F}}_{q}$-basis of $W_{i_{\gamma}}$. If $\{\beta^{(i_{\gamma})}_{1},\ldots,\beta^{(i_{\gamma})}_{R}\}$ is the ${\mathbb{F}}_{q}$-basis of $W_{i_{\gamma}}$, then we define the set of column indexes $\Gamma(\gamma)\subset\{1,\ldots,\ell R\}$ to be the set of indexes of the columns associated to $f(\beta^{(i_{\gamma})}_{1}),\ldots,f(\beta^{(i_{\gamma})}_{R})$. It is obvious that $\gamma\in\Gamma(\gamma)$ and $|\Gamma(\gamma)|=\dim_{{\mathbb{F}}_{q}}(W_{i_{\gamma}})=R=r+\delta-1$, satisfying the first two conditions of Definition 2.21.
To check the third, consider the local code $\operatorname{enc}(\mathcal{M})|_{\Gamma(\gamma)}$ associated with the $\gamma$-th column. For a fixed $1\leq i_{\gamma}\leq\ell$ and basis $\{\beta_{j}^{(i_{\gamma})}\}_{1\leq j\leq R}\subseteq{\mathbb{F}}_{q^{tm}}$ of $W_{i_{\gamma}}=\phi[T-a_{i_{\gamma}}]$, the local code associated with the $\gamma$-th column is as follows:
$\operatorname{enc}(\mathcal{M})|_{\Gamma(\gamma)}=\left\{\left(f(\beta^{(i_{\gamma})}_{1}),\ldots,f(\beta^{(i_{\gamma})}_{R})\right)\bigg|f\in\mathcal{M}\right\}.$
However, we can rewrite $f(x)\in{\mathbb{F}}_{q^{tm}}\langle x\rangle$ restricted to $\phi[T-a_{i_{\gamma}}]$ as
(3.1) $f|_{\phi[T-a_{i_{\gamma}}]}(x)=\sum_{k=0}^{s}\left(g_{k}(a_{i_{\gamma}}^{k}x)\right)=\sum_{k=0}^{s}a^{k}_{i_{\gamma}}g_{k}(x).$
where $g_{k}(x)\in{\mathbb{F}}_{q^{tm}}\langle x\rangle$ with $q$-degree at most r1r-1. If ${\mathbb{F}}_{q^{tm}}\langle x\rangle_{\leq r-1}$ denotes the set of $q$-linearized polynomials with coefficients in ${\mathbb{F}}_{q^{tm}}$ and $q$-degree at most r1r-1, then we have that $f|_{\phi[T-a_{i_{\gamma}}]}(x)$ lives in ${\mathbb{F}}_{q^{tm}}\langle x\rangle_{\leq r-1}$. In light of this, we observe that
$\operatorname{enc}(\mathcal{M})|_{\Gamma(\gamma)}=\left\{\left(p(\beta^{(i_{\gamma})}_{1}),\ldots,p(\beta^{(i_{\gamma})}_{R})\right)\bigg|p\in{\mathbb{F}}_{q^{tm}}\langle x\rangle_{\leq r-1}\right\}=:G.$
By Definition 2.20, $G$ lives in an $(R,r)$-Gabidulin code. Since Gabidulin codes are MRD and using Proposition 2.18, $d_{R}(\operatorname{enc}(\mathcal{M})|_{\Gamma(\gamma)})=R-r+1=\delta$. This proves that $\operatorname{enc}(\mathcal{M})$ is an $\left(tm\times\ell R,tm(s+1)r\right)$ rank-metric code with $(r,\delta)$-rank-locality. In view of this and by Proposition 2.22, we conclude that
Finally, observe that $\operatorname{enc}(\mathcal{M})$ is the set of evaluations of $q$-polynomials of the form $\sum^{s}_{k=0}g_{k}(x)\circ\phi_{T}^{k}(x)$ with $q$-degree at most Rs+r1Rs+r-1 (maximum $q$-degree of $g_{k}(x)$ and $\phi_{T}^{k}(x)$ are at most r1r-1 and $Rs$ respectively) over an ${\mathbb{F}}_{q}$-linearly independent set of size $\ell R\leq tm$ which is $\{\beta_{j}^{(i)}\}_{\begin{subarray}{c}1\leq i\leq\ell\\ 1\leq j\leq R\end{subarray}}$. In other words, $\operatorname{enc}(\mathcal{M})$ is a subcode of an $(\ell R,Rs+r)$ Gabidulin code. Hence
In conclusion, $\operatorname{enc}(\mathcal{M})$ is an optimal $\left(tm\times\ell R,tm(s+1)r,\ell R-Rs-r+1,r,\delta\right)$ matrix rank-metric code. ∎
4. Supersingular Drinfeld Modules
In this section, we show how to realize the ambient space $K$ of our code by choosing proper types of Drinfeld modules. Our goal is to describe enough conditions for the inclusion ${\mathbb{F}}_{q^{m}}(\phi[h])\subseteq K$ to occur, which is critical to be able to work over extension fields whose degree can be controlled.
4.1. Supersingular Drinfeld modules of rank $R>1$
In this subsection we focus on how to determine the ambient space of our code using supersingular Drinfeld modules of rank $R>1$. We start outlining the following result.
Theorem 4.1.
[…]
Proof.
By [19, Excercise 4.2.2], we know that $\tau^{mR}=c\phi_{\mathfrak{p}}$. In other words, $\tau^{mR}=\phi_{\mathfrak{p}c}$. Taking into account the action on $\phi[h]$, we have $\tau^{mR}|_{\phi[h]}=\phi_{\mathfrak{p}c\pmod{h}}|_{\phi[h]}$. Hence, $\tau^{mRb}|_{\phi[h]}=1|_{\phi[h]}$ and $(\tau^{mRb}-1)|_{\phi[h]}=0|_{\phi[h]}$. In other words, we have that $\tau^{Rmb}$ acts like the identity on $\phi[h]$ which forces $\mathbb{F}_{q^{m}}(\phi[h])\subseteq\mathbb{F}_{q^{mRb}}$. ∎
As a particular case of the last result, we show also the following.
@@ 4.1. Supersingular Drinfeld modules of rank $R>1$
Let $\ell<q$, and consider $h=\prod_{i=1}^{\ell}(T-a_{i})\in{\mathbb{F}}_{q}[T]$ where $a_{i}\in\mathbb{F}^{\times}_{q}$ for every $i\in\{1,\ldots,\ell\}$. Let $\mathfrak{p}\in\mathbb{F}_{q}[T]$ be a monic irreducible polynomial of degree $m$, and $\alpha\in\mathbb{F}_{q^{m}}$ be one of its roots. Suppose $R$ is a prime such that $R\nmid m$, and consider the Drinfeld module $\phi:\mathbb{F}_{q}[T]\rightarrow\mathbb{F}_{q^{m}}\{\tau\}$ such that $\phi_{T}=\alpha+\tau^{R}$. Given $m<R\ell$, we have $\mathbb{F}_{q^{m}}(\phi[h])=\mathbb{F}_{q^{mR}}$ if and only if $\mathfrak{p}\equiv 1\pmod{h}$.
@@
By Example 2.12, the Drinfeld module $\phi$ is indeed supersingular. Notice that $\phi[h]=q^{\ell R}$. If $m<\ell R$, then, $\mathbb{F}_{q^{m}}(\phi[h])\supsetneq\mathbb{F}_{q^{m}}$. In light of Theorem 4.1, when $\mathfrak{p}\equiv 1\pmod{h}$, it follows that $\mathbb{F}_{q^{m}}(\phi[h])\subseteq\mathbb{F}_{q^{mR}}$. Since $R$ is a prime, it follows that $\mathbb{F}_{q^{m}}(\phi[h])=\mathbb{F}_{q^{mR}}$. Conversely, if $\mathbb{F}_{q^{mR}}$ is the $h$-division field of $\phi$ then $\phi_{\mathfrak{p}\pmod{h}}|_{\phi[h]}=\tau^{mR}|_{\phi[h]}=1|_{\phi[h]}$ which implies that $\mathfrak{p}\equiv 1\pmod{h}$. ∎
We refer the reader to Section 5.2 for an example of a code that is constructed based on the previous discussion on supersingular Drinfeld modules.
4.2. Drinfeld modules of rank $R=1$
As shown in Section 3, the rank $R$ of the Drinfeld module determines the size of the locality set of the code. This means that using directly rank one Drinfeld modules gives locality sets of size 1, and so it is not possible to recover any information. Nevertheless, if we use powers of Drinfeld modules of rank 1, the construction is still possible.
Lemma 4.3.
[…]